solve the equation by method of variation of parameter?g eye ya e Solve the equation bg method of Variation of Parameter? XS REDMI NOTE 8 PRO AI QUAD CAMERA
Answer #1

Solution 7 ell Solve by Method Veuiation of by using of Parameroes, y-ay + y = x its is in the General Solution form of yW2= ez Some (ez -0 ez 25 we get y yi= er Ya = xer (exy2 ) 2 W.= -(0%)? 24 تركيا 25 # Yo= uigi+ Ugya where ü = wi -2337224 1